OpenStudy (anonymous):

Matt plans to put concrete on a rectangular portion of his driveway. The portion is 8 feet long and 4 inches high. The price of concrete is $98 per cubic yard. The total cost of the concrete Matt needs is $58.07. Which of the following is closest to the width of the portion of the driveway on which Matt plans to put concrete? [1 foot = 12 inches; 1 yard = 3 feet]

OpenStudy (anonymous):

First determine how many cubic yards Matt bought.

OpenStudy (anonymous):

I cant really understand how to solve it. I tried several times but nothing worked

OpenStudy (anonymous):

All the time, i am getting confused

OpenStudy (anonymous):

Ok if he spent $ 58.07 and the price of concrete is $98.00 per cubic yard. Hw many cubic yards did he buy?

OpenStudy (anonymous):

do we need to subtract those two value to find it

OpenStudy (anonymous):

?

OpenStudy (anonymous):

It is a ratio....\[\frac{ 98 }{ 1 } = \frac{ 58.07 }{ x }\] You need to cross multiply to solve for x. Are you familiar with this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok solve for x and let me know.

OpenStudy (anonymous):

98x = 58.07

OpenStudy (anonymous):

98x/98 = 58.07/98 = 0.5925510204082

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

OK so he bought .5925 cubic yards. Looking at the equation we are probably going to have to convert this to cubic inches. So you many cubic feet in a cubic yard?

OpenStudy (anonymous):

3 feet

OpenStudy (anonymous):

1 yard = 3 feet

OpenStudy (anonymous):

Think about a cube that is one yard high, one yard wide and one yard long. So what be in cubic feet?

OpenStudy (anonymous):

8 feet long and 4 inches high

OpenStudy (anonymous):

is it this one your saying?

OpenStudy (anonymous):

I am confused here bit :(

OpenStudy (anonymous):

is it 27 cubic inches?

OpenStudy (anonymous):

Volume of a cube is Length x width x height. If a cube is 1 cubic yard it is Length =1 yard and its width = 1 yard and its height = 1 yard. Thus we get 1 x 1 x1 = 1. Now we need to covert to inches. If there are 36 inches in a yard, we say the length = 36 inches and the width = 36 inches and the height = 36 inches. So a cubic yard in inches would be 36 x 36 x 36 = ? Does this help?

OpenStudy (anonymous):

Not quite

OpenStudy (anonymous):

did you mean 8 inches to multiply three times?

OpenStudy (anonymous):

sorry 4 inches

OpenStudy (anonymous):

like 4*4*4 = 64 inches

OpenStudy (anonymous):

There you go. Now a cubic yard is how many inches on both sides. Don't use numbers from your question, ask youself how many inches in a yard?

OpenStudy (anonymous):

both side I think 64 inches. As you said 1 yard = 27. But I dont know how to convert inches to yard

OpenStudy (anonymous):

There are 36 inches in a yard. So a cubic yard is 36 x 36 x 36 which equals ???

OpenStudy (anonymous):

46656

OpenStudy (anonymous):

Great!!! Now earlier you figured he bought .5926 cubic yards. If 46656 cubic inches = 1 cubic yard. How many cubic inches did he buy? \[\frac{ 1 }{ 46656 }=\frac{ .5926 }{ ? }\]

OpenStudy (anonymous):

x

OpenStudy (anonymous):

am I right?

OpenStudy (anonymous):

No 36

OpenStudy (anonymous):

Cross multiply like you did above.

OpenStudy (anonymous):

isnt it?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So we now know Matt bought 27648 cubic inches of concrete. Now we need to develop a formula. The formula for finding volume is length x width x height = volume. So let's plug in the numbers. What is the length in inches of the driveway.

OpenStudy (anonymous):

so we multiply with 4 inches and 8 feet?

OpenStudy (anonymous):

Close..You have two of the three numbers you need. You do not know the width that's what we are trying to solve. What is the length of the 8 foot long driveway in inches?

OpenStudy (anonymous):

4 inches

OpenStudy (anonymous):

8 feet long = 4 inches????

OpenStudy (anonymous):

I am really sorry trouble, I am really struggling with it. That's why, it is taking long

OpenStudy (anonymous):

Not a problem....8 feet equals 96 inches.. 12 inches in a foot x 8 = 96. So we have one number of our formula 96 x height x width = Volume. Looking at the question what is given as the height of the driveway?

OpenStudy (anonymous):

4 inches high

OpenStudy (anonymous):

Great!!!! Now we have two numbers of the formula 96 x 4 x width = Volume. Now they want us to solve for width so that is not given. Earlier we figured the volume as 27648. So our formula looks like this 96 x 4 x width = 27648. Now you need to solve for the width. Can you do that or do I need to help?

OpenStudy (anonymous):

I am not sure, would mind help me here too? Can you tell me how do find 8 feet = 96 inch?

OpenStudy (anonymous):

There are how many inches in 1 foot?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

If there are 8 feet and there are 12 inches in a foot. Then 8 (number of feet) x 12(inches in a foot) = 96

OpenStudy (anonymous):

Oh I get now how you got 96

OpenStudy (anonymous):

Back to the original problem. Our solution looks like this 96 x 4 x width = 27648. We need to isolate the unknown (width) on one side of the equation. So let's start out by multiplying 96 x 4 . What do you get?

OpenStudy (anonymous):

384

OpenStudy (anonymous):

So our equation now looks like this 384 x width = 27648. Now can you solve for the width?

OpenStudy (anonymous):

I am trying :)

OpenStudy (anonymous):

I know you are

OpenStudy (anonymous):

72

OpenStudy (anonymous):

Winner winner chicken dinner... Excellent. We now know the width is 72 inches...how many feet is that???

OpenStudy (anonymous):

72*1/12 = 6 feet. Thanks so so much for teaching me so hard!!! :) I was having really hard time to understand it

OpenStudy (anonymous):

You did great and it was a pleasure... :-}

OpenStudy (anonymous):

It is all because of you, entire credit goes to you :) thanks a lot!!!